OpenStudy (anonymous):

Determine whether the given point is a solution of 3x+y=4 1. (0,4) 2. (1,-1)

OpenStudy (anonymous):

plug in (x=0,y=4) into your equation and see if it equals 4. if it doesnt, its not a solution. do the same for part 2

OpenStudy (anonymous):

let's see

OpenStudy (anonymous):

plug in the points

OpenStudy (anonymous):

Can you help me with the show work and answering the question?

OpenStudy (anonymous):

3(0) + (4) = 4

OpenStudy (anonymous):

4 = 4

OpenStudy (anonymous):

3(0)+4=4 4=4 so its a solution

OpenStudy (anonymous):

So, 1 is correct

OpenStudy (anonymous):

3(1) +(-1) = 4

OpenStudy (anonymous):

3-1 = 4

OpenStudy (anonymous):

since (0,4) is (x,y). you can do the same for second part. the second part is not a solution, can u say why?

OpenStudy (anonymous):

2 =4 this is not correct. SO, 2 is not correct

OpenStudy (anonymous):

Only number 1 works as a solution

OpenStudy (anonymous):

or faisalalif can give you the answer haha

OpenStudy (anonymous):

(0,4) only work lol

OpenStudy (anonymous):

this is easy stuff @Flop

OpenStudy (anonymous):

its easy for us, not her

OpenStudy (anonymous):

i think this is Pre - Algebra